OpenStudy (anonymous):

Find the Inverse y=7x^2−3

OpenStudy (decentnabeel):

\[\mathrm{Inverse\:of}\:7x^2-3:\quad -\frac{\sqrt{x+3}}{\sqrt{7}},\:\frac{\sqrt{x+3}}{\sqrt{7}}\]